OpenStudy (anonymous):

Simplify the exponential expression: (2/5)^-3?

OpenStudy (anonymous):

\[(\frac{ 2 }{ 5 })^(-3) = (\frac{ 5 }{ 2 })^(3) = \frac{ 125 }{ 8 } \]

OpenStudy (anonymous):

-3 and 3 are in powers of 2/5 and 5/2

OpenStudy (anonymous):

thnks